The time to failure T of a component has probability density f ( t ) as shown
(b) Derive the corresponding survivor function R ( t ) .
(c) Derive the corresponding failure rate function z ( t ) , and make a sketch of z(t)

Note: The f(t) is a valid pdf (so we can obtain c or the height of the triangle). Information are enough to solve this problem.

f(t) a -b a b Time t Fig. 2.27 Probability density (Problem 2.28)
Answer #1

  0 - (a+b)E la, a + d 0 2 a-(a+b)

Rightarrow f(t)=left{egin{matrix} ;;rac{c}{b}[t-a+b] &;;;tin [a-b,a] -rac{c}{b}[t-a-b] &;;;tin [a,a+b] end{matrix} ight.

Now   f(t)dt 1 i.e. area under the curve of f(t) = 1 Rightarrow (1/2)[(a+b)-(a-b)]c = 1 Rightarrow bc = 1 Rightarrow c = 1/b

. f(t) (t-a-b) te[, a+b

f (t)dt

b2 Ja-b b2 Ja-b

a,a + b 2 2b2

  (t - ab)2 E a,a+ b

(b)

018102- R(t) 1- F(t)11 426 E a, a + b 2 2b2

(c)

(t) = f(t) t-atb 2(t-a-b) h t E [a, a + b

tela-b,a h(t)